OpenStudy (anonymous):

I'm having trouble working this problem out. Help pls? Find dy/dx by implicit differentiation. (x − y)/(4x + 10y) = 8x

OpenStudy (inkyvoyd):

1/4-y/4x+10y=8x

OpenStudy (inkyvoyd):

Can you get it from there?

OpenStudy (anonymous):

I'm new at this by the way. lol. so, are you dividing the function by 1/4-y or how did you get that part?

OpenStudy (anonymous):

I understand derivatives, but the differentiation is throwing me off..

OpenStudy (inkyvoyd):

I just split that fraction up,

OpenStudy (inkyvoyd):

(x − y)/4x + 10y = 8x, assuming that the 4x is in parenthesis\[(x-y)/(4x)+10y=8x, x/(4x) -y/(4x)+10y=8x\]

OpenStudy (inkyvoyd):

Now, multiply both sides by 4x, and tell me if you hav trouble

OpenStudy (anonymous):

so multiplying by 4x on each side would eliminate the 4x on the denominator for only x, and not y?

OpenStudy (anonymous):

from my notes, I'm instructed to use the quotient rule, however if this method is easier and effective I'll use your way.

OpenStudy (anonymous):

maybe I didn't correctly display the function. Does this make it a different problem? 

OpenStudy (inkyvoyd):

Yes, it does. Actually I know what you are supposed to do; let me show you.,

OpenStudy (anonymous):